40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6053     Entry time: Wed Nov 30 13:52:09 2011
Author: Zach 
Type: Update 
Category: RF System 
Subject: EOM temp stabilization ineffectual 

This recent off/on run proved what I was afraid of: the temperature stabilization setup appears to do nothing except very near DC. The plot that Kiwamu posted is misleading because the "uncontrolled" data stretch at the beginning actually had the heater injecting random noise (since the circuit was broken). Below are some plots (sorry in advance for their crappiness---the plot tools wouldn't let me print to file for some reason):

Time series of the temp monitor, the heater monitor, and the 11- and 55-MHz RFAM monitors. The heater was disconnected at ~2:20 UTC, and the temperature is seen to equalize to the surroundings (note that the TEMP_MON is inverted, so positive change means decreasing temperature). The heater was reconnected by Kiwamu around 10:40 UTC, and you can see the temperature being driven back to the zero point by the loop. Note that the temperature was still decreasing at a fair rate when the heater is re-engaged---this could mean that we really need to take longer samples.

Screen_shot_2011-11-30_at_1.31.44_PM.png

 

Spectra and coherence of the 11- and 55-MHz RFAM monitors before and after the control was re-engaged. It appears that the 11-MHz signal is unaffected while the 55-MHz signal actually gets worse. This also seems evident from the noisiness in the time series for that signal above (top right). Bad, bad, bad. 

Screen_shot_2011-11-30_at_1.29.48_PM.png

 

Spectrum of the EOM temperature signal before and after control was re-engaged. There seems to be no change whatsoever. Of course, as mentioned before, this signal seems to be close to the digitization noise level as seen in DV. I am not sure what the ADC noise looks like at these low frequencies. In case someone knows, the magnitude of this signal in counts is ~0.1 ct/rHz at 10 mHz; is this too low? Another thing to note is that the noise level in K is pretty low! I would be surprised if the RTD can really see 10 uK/rHz at and below 10 mHz.

 

Screen_shot_2011-11-30_at_1.21.16_PM.png

 

I need to try and increase the gain of the servo to see if I can get it much higher without it becoming unstable.

ELOG V3.1.3-